OpenStudy (anonymous):

The median of a set of data is the middle value. In order to find the median you must first arrange the values in order from smallest to largest. Then find the middle value. Below is a listing of lengths of boards that can be used. What is the medan value? 5 1/2 ft. 5 7/8 ft. 5 2/20 ft. 5 5/6 ft. 5 4/12 ft. A. What will you need to do first before you can arrange the lengths in order from smallest to largest? How will this be accomplished? Write your answers to the questions in sentences. B. Show all work to arrange the values in order. *Continued Below*

OpenStudy (anonymous):

C. Write the original values in order from smallest to largest. D. What is the median value

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You have a choice. 1) Make all the fractions have the same denominator. 2) Use the decimal expansion of each fraction. 3) Magic. Quite a few students try method #3. I don't recommend it.

OpenStudy (anonymous):

OK. @tkhunny I can do the method #1.

OpenStudy (tkhunny):

Method #1 can be a little annoying for a large collection of inconvenient numbers, but that doesn't mean it won't work. Let's see it.

OpenStudy (anonymous):

@tkhunny 120

OpenStudy (anonymous):

could you help @alakara_kibo

OpenStudy (anonymous):

sure givve me a moment :)

OpenStudy (anonymous):

Please answer ALL parts @alakara_kibo

OpenStudy (anonymous):

I would put the fractions in decimal form

OpenStudy (anonymous):

5 1/2 = 5.50 5 7/8 = 5.875 5 2/20 = 5.10 5 5/6 = 5.83 5 4/12 = 5.33 does that help ?

OpenStudy (anonymous):

@kelliegirl33 Sort of thank you but I'm kind of looking for the answer for all parts.

OpenStudy (anonymous):

A) find the decimal form, and this will be accomplish it by finding the decimal , B)5.2 ,5. 33, 5 .875 , 5.83 ,5.5. C) 5 2/20, 5 4/12, 5 7/8 ,5 5/6, 5 1/2. D) the value is 5 7/8

OpenStudy (anonymous):

there u go HBO102

OpenStudy (anonymous):

Look at there very first comment. It has parts C and D. Could you answer them? @alakara_kibo

OpenStudy (anonymous):

i did it caps C and D

OpenStudy (anonymous):

put them in order from smallest to largest and the median is the middle number.. 5 2/20 , 5 4/12, 5 1/2, 5 5/6, 5 7/8....median is 5 1/2 Before you can arrange from smallest to largest, you either have to turn them into fractions with a common denominator, or turn them into decimals.

OpenStudy (anonymous):

ok that is fine out of all of them i missed that one sorry

OpenStudy (anonymous):

@kelliegirl33 Thank you so very much !! I might use the common denominator as that's what we used in this chapter but I appreciate you taking your time :) Thank you @alakara_kibo for taking your time to help me as well :)

OpenStudy (anonymous):

you need help with the common denominator ?

OpenStudy (anonymous):

ur welcome:)

OpenStudy (anonymous):

I should be good @kelliegirl33. I know the common denominator is 120 so it is kinda easy now. Thanks.

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

Actually I might need help @kelliegirl33 ....

OpenStudy (anonymous):

I have the 120 slapped onto each fraction. But I'm having trouble putting them from smallest to largest.

OpenStudy (anonymous):

Could you help @alakara_kibo ?? Read the comment above ^

OpenStudy (anonymous):

give me a few minutes ok ty :)

OpenStudy (anonymous):

@kelliegirl33 I need help putting them into smallest and largest with the common denominator of 120.

OpenStudy (anonymous):

5 1/2 = 11/2 5 7/8 = 47/8 5 2/20 = 5 1/10 = 51/10 5 5/6 = 35/6 5 4/12 = 5 1/3 = 16/3 so we have 11/2, 47/8, 51/10, 35/6, 16/3 your right...least common denominator is 120 11/2 = 660/120 47/8 = 705/120 51/10 = 612/120 35/6 = 700/120 16/3 = 640/120 again, in order from smallest to largest... 5 2/20 = 51/10 = 612/120 5 4/12 = 16/3 = 640/120 5 1/2 = 11/2 = 660/120 5 5/6 = 35/6 = 700/120 5 7/8 = 47/8 = 705/120

OpenStudy (anonymous):

OK. Thank you sooo much !! I really appreciate it :) @kelliegirl33

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

by the way @kelliegirl33 thanks for a little lessons on fractions lol

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)